Mark and Sam formed a management consulting partnership on January 1, 2009. The fair value of the net assets invested by each partner follows:

                                          Mark            Sam

Cash                            $10,000     $20,000

Accounts receivable      10,000         5,000

Office equipment           40,000               -

Land                                -             50,000

Accounts payable        15,000         12,000

During the year, Mark withdrew $10,000 and Sam withdrew $16,000. Net profit for 2009 was $30,000, which is to be allocated based on the original net capital investment.

Requirement:                                                                           

1-   Prepare the appropriate journal entries to record the initial investment and drawings in the partnership for both partners.

2-   Prepare closing entries.

3-   Calculate the ending balance for capital for both partners.



1
Expert's answer
2021-11-07T19:40:04-0500

Let's present the data in EXCEL to calculate the percentage for profit distribution:



1.Debit cash credit settlements with founders - 10 000

Debit cash credit settlements with founders - 20 000

Debit Accounts receivable credit settlements with founders - 10 000

Debit Accounts receivable credit settlements with founders - 5 000

Debit Office equipment  credit settlements with founders - 40 000

Debit Land  credit settlements with founders - 50 000

Debit settlements with founders credit Accounts payable - 15 000

Debit settlements with founders credit Accounts payable - 12 000

Debit settlements with founders credit Authorized capital - 45 000

Debit settlements with founders credit Authorized capital - 63 000

"%Mark=\frac{45 000}{45000+63000}=0.416" "%Mark=\frac{45000}{45000+63000}=0.416" "Mark=\frac{45000}{45000+63000}=0.416" or 42%

"Sam=\frac{63000}{45000+63000}=0.583" or 58%

Debit settlements with founders credit cash - 10 000

Debit settlements with founders credit cash - 16 000

"profit Mark=0.42\times30000=12600"

"profit Sam=0.58\times30000=17400"

2.3.Let's imagine the balances taking into account the transactions